margo and tony had a running race. margo ran 10 feet farther than tony. together they ran 64 feet. how far did each person run?
Answer by checkley71(8403)

(T+10)+T=64
2T=64-10
2T=54
T=54/2
T=27 FEET FOR TONY.
MARGE=27+10=37
PROOF
27+37=64
64=64